How long will it take for any amount of money to triple if invested at 6% annual interest and compounded quarterly and compounded monthly
Answer by mananth(12270)

Either you can take x as the principal Or %100
I take $100
Principal P = 100
Amount= 300
years=n
compounded (t) 4
Rate = 6 0.02
Amount = P*((n+r)/n)^n

300 = 100 *( 1 + 0.02 )^ n*t
3 = *( 1 + 0.02 )^ n* 4
ln 3 = 4 n *ln 1.02
1.1 = 4 0.01 *n
18.45 = n
18.45 years
-----------------
Monthly
Principal P = 100
Amount= 300
years=n
compounded (t) 12
Rate = 6 0.01
Amount = P*((n+r)/n)^n

300 = 100 *( 1 + 0.01 )^ n*t
3 = *( 1 + 0.01 )^ n* 12
ln 3 = 12 n *ln 1.01
1.1 = 12 0 *n
18.36 = n
m.ananth@hotmail.ca